OpenStudy (anonymous):

The circle formed by the blades of a fan has a circumference of eighty-one and five-sevenths inches. What is its diameter? Use twenty-two sevenths for pi. twelve inches eighteen inches twenty-one inches twenty-six inches

OpenStudy (anonymous):

@ganymede

OpenStudy (anonymous):

diameter * pi = cicumference

OpenStudy (anonymous):

\[\frac{ circumference }{ \pi } = diameter\]

OpenStudy (anonymous):

81 * 7 = 567, 567+5 = 572, so 81 and 5/7 can be re-written 572/7

OpenStudy (anonymous):

\[\frac{ \frac{ 572 }{7} }{ \frac{ 22 }{ 7}}\]

OpenStudy (anonymous):

so 26 in.

OpenStudy (anonymous):

yep